# BPMN Assistant: An LLM-Based Approach to Business Process Modeling

### v1

@@ Abstract
This paper presents BPMN Assistant, a tool that leverages Large Language Models (LLMs) for natural language-based creation and editing of BPMN diagrams. A specialized JSON-based representation is introduced as a structured alternative to the direct handling of XML to enhance the accuracy of process modifications. Process generation quality is evaluated using Graph Edit Distance (GED) and Relative Graph Edit Distance (RGED), while editing performance is evaluated with a binary success metric. Results show that JSON and XML achieve similar similarity scores in generation, but JSON offers greater reliability, faster processing, and significantly higher editing success rates. We discuss key trade-offs, limitations, and future improvements. The implementation is available at https://github.com/jtlicardo/bpmn-assistant.
@@ 1 Introduction
Business Process Model and Notation (BPMN) has long served as a standard for modeling business processes, enabling organizations to visualize and optimize their workflows. However, the complexity inherent in creating, editing, and interpreting BPMN diagrams poses significant challenges, particularly for individuals without specialized training [1]. As Recker [2] demonstrates through extensive research, BPMN’s over-engineered nature and the prevalent lack of formal training among its users create substantial barriers to effective adoption. This bottleneck often results in inefficiencies and a reliance on experts, which can increase operational costs and delay decision-making.
A significant challenge in modern organizations is the communication gap between IT departments and business stakeholders. While IT professionals are comfortable with formal modeling notations and technical specifications, business users typically express processes in natural language and informal descriptions. A systematic review by [3] demonstrates that this disconnect leads to misunderstandings, requirements misalignment, and implementation delays, with communication barriers persisting despite significant investments in alignment initiatives. Furthermore, valuable process knowledge frequently remains trapped in the minds of domain experts or scattered across various informal documents, making it difficult to capture and formalize this expertise [4].
@@
This work presents BPMN Assistant, a system that leverages Large Language Models to address the fundamental challenges in business process modeling. While recent research has demonstrated the potential of AI-assisted process modeling, the critical aspect of model modification and maintenance remains largely unexplored. Our work advances the state of the art by demonstrating the effectiveness of LLMs in editing existing BPMN diagrams through natural language instructions.
Our evaluation reveals significant improvements in editing success rates compared to baseline approaches, particularly when using structured representations for model manipulation. These results suggest that LLMs can effectively bridge the gap between informal process descriptions and formal models, offering a promising solution for maintaining and updating business process documentation.
@@ 2 Related Work
The integration of Large Language Models (LLMs) into Business Process Management (BPM) has led to significant developments in process automation and modeling. The introduction of specialized tools like the BPMN-Chatbot [19] and ProMoAI [20] has demonstrated the practical applications of this technology. Initial explorations by Berti and Qafari [21] into using GPT-4 and Bard (Gemini) for process mining tasks showed promising results in interpreting both procedural and declarative models, highlighting the potential of LLMs in the BPM domain. These developments have led to comprehensive frameworks for process modeling with LLMs [22], showing promising results in both automation and model quality.
@@
ProMoAI, introduced by Kourani et al. [20], represents a significant development in LLM-driven process modeling. While the tool incorporates prompt engineering and error handling techniques, it currently generates Graphviz graph visualizations rather than process models based on standardized BPMN notation. The tool’s effectiveness was thoroughly evaluated in a subsequent study [26], which demonstrated its capabilities when paired with different LLMs, with Claude 3.5 Sonnet emerging as the best-performing model for generating high-quality process models. Despite using non-standard notation, ProMoAI’s approach demonstrates strong potential for using LLMs to assist in process visualization tasks.
@@
While BPMN-Chatbot demonstrates notable strengths, particularly in token efficiency and correctness rates, its evaluation primarily focuses on the generation of process models. The editing capabilities, which are essential for iterative refinement and interactive process modeling, have not been rigorously evaluated. Addressing this gap could further validate its utility in real-world scenarios where users frequently modify and adapt process models.
@@ 3 System Architecture
The BPMN Assistant system architecture, as illustrated in Figure 1, is designed to facilitate seamless interaction between users and BPMN models through natural language inputs. The system is composed of three primary components: the Python-based backend, the BPMN layout server, and the Vue.js frontend. Each component plays a critical role in ensuring the system’s functionality, efficiency, and usability.
@@ 3.1 Backend
The backend, implemented in Python, serves as the core computational engine of the system. Its primary responsibilities include handling user inputs, interacting with the Large Language Model (LLM), managing BPMN diagrams, and providing APIs for the frontend. The backend leverages the FastAPI framework, chosen for its performance and simplicity in developing RESTful APIs.
The backend serves as the first point of contact for user inputs, analyzing these inputs to determine the user’s intent. This can include conversational queries, such as asking questions about BPMN concepts, or operational commands, such as creating or modifying BPMN diagrams. When a conversational intent is recognized, the backend communicates with the LLM to generate a natural language response. If the input requires an operational response, the backend constructs a prompt to instruct the LLM to generate or modify a BPMN diagram. These diagrams are initially represented in JSON format and subsequently converted to BPMN XML for compatibility with visualization tools. Additionally, the backend supports the uploading of BPMN files, enabling users to query or modify existing diagrams. The backend includes robust error handling and validation mechanisms to ensure that outputs from the LLM are accurate and reliable. If invalid JSON is generated, the backend retries the process or notifies the user of the error. By integrating with the BPMN layout server, the backend also ensures that diagrams are enriched with graphical information for enhanced visualization.
@@
Our system integrates multiple models from different providers, each chosen for its specific strengths and use cases. Table 1 provides an overview of the supported models and their capabilities.
@@
OpenAI GPT-4o Multimodal model capable of processing text, audio, image, and video [29].
GPT-4o mini Fast and affordable model optimized for simpler tasks
o3-mini Cost-efficient reasoning model for coding and science use-cases.
Anthropic Claude 3.5 Sonnet Advanced model featuring a 200K token context window for enhanced reasoning.
Google Gemini 2.0 Flash Multimodal model with with low latency and a 1M context window.
Fireworks AI Llama 3.3 70B Meta’s latest open-source LLM [30].
@@
Model Input Output
o3-mini 1.10 4.40
GPT-4o 2.50 10.00
GPT-4o mini 0.15 0.60
Claude 3.5 Sonnet 3.00 15.00
Gemini 2.0 Flash 0.10 0.40
Qwen 2.5 72B 0.90 0.90
Llama 3.3 70B 0.90 0.90
Deepseek V3 0.75 3.00
Table 2: Pricing for AI Models (in USD per million tokens)
This diverse selection ensures flexibility and performance optimization across different AI-driven tasks.
@@
The backend facilitates operations on a variety of BPMN elements, including tasks, gateways, and events. These supported elements, detailed in Table 3, enable the system to cover a comprehensive range of BPMN functionalities, enhancing its applicability to diverse workflows.
Category Elements
Tasks General tasks, user tasks, service tasks
Gateways Exclusive gateway, parallel gateway
Events Start event, end event
Table 3: Supported BPMN Elements
@@ 3.2 BPMN Layout Server
The BPMN layout server, implemented using Node.js with the Express.js framework, is dedicated to augmenting BPMN diagrams with graphical information. This server employs the bpmn-auto-layout npm library to generate DI (Diagram Interchange) information, which includes the graphical coordinates of BPMN elements.
@@ 3.4 Data Flow
The system processes data through a series of well-defined steps to ensure efficient operation and user satisfaction. Initially, user inputs are submitted via the frontend and transmitted to the backend, where intent recognition determines the appropriate response. If the input is conversational, the backend queries the LLM to generate a natural language reply. For operational inputs, the backend prompts the LLM to produce BPMN diagrams in JSON format. This JSON is then converted into BPMN XML and sent to the BPMN layout server. The layout server enriches the XML with graphical information and returns it to the backend. Finally, the enriched BPMN diagram is sent to the frontend, where it is displayed in the BPMN canvas. Users can choose to continue interacting with the diagram or download it for their records.
@@ 3.5 JSON Representation of BPMN Diagrams
The BPMN JSON representation uses a sequence of elements to describe the process. Each element is executed in order based on its position in the process array unless gateways (exclusive or parallel) specify branching paths. The following describes the structure of various BPMN elements supported by the system.
@@
Tasks represent activities within a process and are specified with the type field. The system supports three task types: general tasks, user tasks, and service tasks. When defining tasks, it is recommended to specify the specific task type (userTask or serviceTask) when possible.
    "type": "task" | "userTask" | "serviceTask",
    "id": String,
    "label": String // short task description
@@
Events mark the beginning or end of a process. The system supports two event types: startEvent and endEvent. These events are critical for defining the boundaries of a workflow.
    "type": "startEvent" | "endEvent",
    "id": String,
    "label": String // optional event label
@@
Gateways determine the flow of a process by introducing branching or parallel paths. Two types of gateways are supported: exclusive and parallel gateways.
Exclusive gateways allow for conditional branching based on specified conditions. Each branch contains a condition and an optional path array of elements executed if the condition is met.
    "type": "exclusiveGateway",
    "id": String,
    "label": String, // gateway label
    "has_join": true | false, // whether the gateway merges branches
@@
            "condition": String, // condition for the branch
            "path": [], // optional sequence of elements
            "next": String // optional next element ID
Parallel gateways execute multiple branches simultaneously. Each branch is represented as an array of elements executed in parallel. Synchronization is automatically handled by the system, eliminating the need to define converging elements explicitly.
@@
        [], // branch 1 elements
        []  // branch 2 elements
@@ 3.6 Process Editing Functions
The system supports a set of specialized functions for modifying BPMN diagrams. These functions, outlined in Table 4, enable precise control over process elements while maintaining the structural integrity of the diagram. When the LLM receives an editing request, it analyzes the natural language input and determines which function(s) to call to achieve the desired modifications.
@@ 4 Evaluation
Our evaluation methodology deliberately focuses on structural and syntactic assessment through graph similarity measures. This approach was selected based on the research objectives of comparing intermediate representations for BPMN generation and modification. It is important to note that this evaluation does not encompass semantic evaluation approaches such as those based on Petri nets or conformance checking from process mining. Similarly, we do not focus on execution performance through simulation studies, nor do we conduct comprehensive usability evaluations. While these aspects represent valuable dimensions for assessing process modeling tools, they fall outside the scope of the current research, which primarily examines the efficacy of different representation approaches in generating structurally accurate BPMN diagrams from natural language descriptions. The selected evaluation metrics allow for direct comparison between JSON and XML approaches while providing insight into the structural correctness of the generated models.

### v2

@@ Abstract
This paper presents BPMN Assistant, a tool that leverages Large Language Models for natural language-based creation and editing of BPMN diagrams. While direct XML generation is common, it is verbose, slow, and prone to syntax errors during complex modifications. We introduce a specialized JSON-based intermediate representation designed to facilitate atomic editing operations through function calling. We evaluate our approach against direct XML manipulation using a suite of state-of-the-art models, including GPT-5.1, Claude 4.5 Sonnet, and DeepSeek V3. Results demonstrate that the JSON-based approach significantly outperforms direct XML in editing tasks, achieving higher or equivalent success rates across all evaluated models. Furthermore, despite requiring more input context, our approach reduces generation latency by approximately 43% and output token count by over 75%, offering a more reliable and responsive solution for interactive process modeling.
@@ 1 Introduction
Business Process Model and Notation (BPMN) [5] has long served as a standard for modeling business processes, enabling organizations to visualize and optimize their workflows. However, the complexity inherent in creating, editing, and interpreting BPMN diagrams poses significant challenges, particularly for individuals without specialized training [37]. As Recker [36] demonstrates through extensive research, BPMN’s over-engineered nature and the prevalent lack of formal training among its users create substantial barriers to effective adoption. This bottleneck often results in inefficiencies and a reliance on experts, which can increase operational costs and delay decision-making.
A significant challenge in modern organizations is the communication gap between IT departments and business stakeholders. While IT professionals are comfortable with formal modeling notations and technical specifications, business users typically express processes in natural language and informal descriptions. A systematic review by Njanka et al. [31] demonstrates that this disconnect leads to misunderstandings, requirements misalignment, and implementation delays, with communication barriers persisting despite significant investments in alignment initiatives. Furthermore, valuable process knowledge frequently remains trapped in the minds of domain experts or scattered across various informal documents, making it difficult to capture and formalize this expertise [10].
@@
This work presents BPMN Assistant, a system that leverages LLMs to address these challenges. While recent research has demonstrated the potential of AI-assisted process modeling, the critical aspect of model modification and maintenance remains largely unexplored. Our work advances the state of the art by demonstrating the effectiveness of LLMs in editing existing BPMN diagrams through natural language instructions. Rather than framing BPMN modeling as a one-shot generation problem, this work is conceptualized around the idea of incremental process transformation. We argue that BPMN models should be treated as mutable process structures that can be reliably modified through a constrained set of well-defined operations, instead of being repeatedly regenerated in their entirety. This conceptual perspective motivates the separation of process logic from BPMN 2.0 XML syntax and forms the basis for the structured intermediate representation proposed in this study.
1.1 Research Objectives
This study addresses the following research questions:
  • RQ1: How does a structured intermediate representation (JSON) compare to direct standard notation (XML) in terms of generation reliability and editing success rates?
  • RQ2: To what extent does a function-based editing approach enable open-weights models to perform complex modeling tasks previously reserved for proprietary frontier models?
  • RQ3: How does the trade-off between increased input context and reduced output complexity impact overall system latency and operational efficiency?
1.2 Contribution
We propose a novel approach that utilizes a structured, JSON-based intermediate representation to abstract away the syntactic complexity of BPMN 2.0 XML. By shifting the generative task from producing verbose XML to manipulating a concise JSON structure, our system allows LLMs to focus on process logic rather than formatting rules. At a conceptual level, this approach reframes BPMN editing as a controlled transformation problem, where modifications are expressed as atomic logical operations rather than as low-level syntactic changes to XML documents.
Our evaluation reveals significant improvements in editing success rates compared to baseline approaches, particularly when using structured representations for model manipulation. Furthermore, we demonstrate that this approach significantly reduces processing latency and output token costs, making it a viable solution for interactive tools.
The implementation of BPMN Assistant is publicly available at https://github.com/jtlicardo/bpmn-assistant.
@@ 2 Related Work
The integration of LLMs into Business Process Management (BPM) has led to significant developments in process automation and modeling. The introduction of specialized tools like the BPMN-Chatbot [20] and ProMoAI [23] has demonstrated the practical applications of this technology. Initial explorations by Berti and Qafari [3] into using GPT-4 and Bard (Gemini) for process mining tasks showed promising results in interpreting both procedural and declarative models, highlighting the potential of LLMs in the BPM domain. These developments have led to comprehensive frameworks for process modeling with LLMs [22], showing promising results in both automation and model quality.
2.1 Process Extraction and Prompting Strategies
@@
2.2 Interactive Modeling Tools
ProMoAI, introduced by Kourani et al. [23], represents a significant development in LLM-driven process modeling. The tool prompts an LLM to generate constrained Python code for constructing an intermediate POWL representation and incorporates prompt-engineering and iterative error-handling mechanisms to improve reliability. The resulting models are then transformed into standard notations and can be viewed and exported as BPMN and Petri nets (PNML).
The framework underpinning ProMoAI was evaluated in a subsequent benchmark study [21], which compared multiple LLMs using conformance checking (harmonic mean of fitness and precision against ground-truth event logs) as a model-quality measure. In this evaluation, Claude 3.5 Sonnet achieved the highest average quality score (0.93), approaching the ground-truth baseline (0.98).
@@ 2.2 Interactive Modeling Tools
While BPMN-Chatbot demonstrates notable strengths, particularly in token efficiency and correctness rates, its empirical evaluation does not isolate the robustness of iterative editing in the feedback loop (e.g., edit success rate over multi-step refinements, structural validity guarantees, or minimal-change behavior), beyond a preliminary technology acceptance test.
Hörner et al. [16] introduce BPMNGen, an LLM-based conversational framework for generating BPMN 2.0 process models from natural-language descriptions and iteratively refining them through user prompts. Their work is distinguished by a comprehensive human-centered evaluation, including controlled user studies assessing semantic alignment, cognitive load, acceptability, and comprehension performance of LLM-generated models compared to expert-created ones. The results indicate that LLM-generated models can achieve expert-level semantic quality for simple and moderately complex processes, while expert intervention remains beneficial for highly complex scenarios.
In contrast, BPMN Assistant focuses on a different aspect of the problem: robust and fine-grained interactive BPMN editing, emphasizing structural correctness, editing reliability, latency, and token efficiency across multiple LLM backends. Rather than evaluating model comprehensibility, our work addresses the engineering challenges of incremental model manipulation and validation, which are complementary to the human-centered quality dimensions studied in BPMNGen.
2.3 Automated Generation and Evaluation Frameworks
@@
Axis ProMoAI [23, 21] BPMN-Chatbot [20] BPMNGen [16] BPMN Assistant
[…]
Output notation BPMN + PNML BPMN BPMN BPMN
Evaluation focus conformance-based model quality correctness + token efficiency + acceptance semantic alignment + cognitive load + acceptability + comprehension GED/RGED structural fidelity + failure rate
Evaluation method automated automated + user study user study automated
Table 1: Comparison of interactive modeling tools by research objective and evaluation emphasis.
@@ 3 System Architecture
The system architecture is designed to operationalize the core conceptual assumption of this work: that reliable BPMN modeling with LLMs requires explicit separation between process semantics, editing logic, and concrete BPMN serialization. The BPMN Assistant system architecture, as illustrated in Figure 1, is designed to facilitate seamless interaction between users and BPMN models through natural language inputs. The system is composed of three primary components: the Python-based backend, the BPMN layout server, and the Vue.js frontend. Each component plays a critical role in ensuring the system’s functionality, efficiency, and usability.
@@ 3.1 Backend
The backend, implemented in Python, serves as the core computational engine of the system. Its primary responsibilities include handling user inputs, interacting with the LLM, managing BPMN diagrams, and providing APIs for the frontend. The backend leverages the FastAPI framework, chosen for its performance and simplicity in developing RESTful APIs.
The backend serves as the first point of contact for user inputs, analyzing these inputs to determine the user’s intent. This can include conversational queries, such as asking questions about BPMN concepts, or operational commands, such as creating or modifying BPMN diagrams. When a conversational intent is recognized, the backend communicates with the LLM to generate a natural language response. If the input requires an operational response, the backend constructs a prompt to instruct the LLM to generate or modify a BPMN diagram. These diagrams are initially represented in JSON format and subsequently converted to BPMN XML for compatibility with visualization tools. Additionally, the backend supports the uploading of BPMN files that conform to the supported BPMN subset, enabling users to query or modify existing diagrams within the constraints of the system’s JSON representation. The backend includes robust error handling and validation mechanisms to ensure that outputs from the LLM are accurate and reliable. If invalid JSON is generated, the backend retries the process or notifies the user of the error. By integrating with the BPMN layout server, the backend also ensures that diagrams are enriched with graphical information for enhanced visualization.
@@
Our system supports the usage of a wide range of LLMs from different providers, each chosen for its specific strengths and use cases. Table 2 provides an overview of the supported models and their capabilities.
@@
OpenAI GPT-5.1 OpenAI’s flagship model for coding and agentic tasks.
GPT-5 mini Faster, more cost-efficient version of GPT-5, used for well-defined tasks.
GPT-4o Earlier-generation multimodal model with broad cross-modal capabilities [32].
Anthropic Claude 3.5 Sonnet Earlier-generation model with strong reasoning and long-context capabilities.
Claude 4.5 Sonnet Frontier model optimized for agentic reasoning and complex coding tasks.
Google Gemini 2.0 Flash Multimodal model with with low latency and a 1M context window. [6]
Fireworks AI Llama 3.3 70B Meta’s open-source text-only LLM [14].
@@
The backend facilitates operations on a wide range of BPMN elements to support diverse modeling scenarios. In the category of tasks, the system supports generic tasks, user tasks, service tasks, send tasks, receive tasks, business rule tasks, manual tasks, and script tasks. To manage process flow logic, the system supports exclusive, parallel, and inclusive gateways.
The system also provides support for events, distinguishing between start, end, and intermediate types. Specifically, it supports generic, timer, and message start events; generic and message end events; as well as generic and message intermediate throw events, and generic, timer, and message intermediate catch events. This selection of elements enables the modeling of specific executable processes and event-driven workflows.
@@ 3.2 BPMN Layout Server
The BPMN layout server, implemented using Node.js with the Express.js framework, is dedicated to augmenting BPMN diagrams with graphical information. This server employs the bpmn-auto-layout [^1] https://github.com/bpmn-io/bpmn-auto-layout npm library to generate DI (Diagram Interchange) information, which includes the graphical coordinates of BPMN elements.
@@ 3.4 Data Flow
The system processes data through a structured sequence of interactions, as illustrated in Figure 1. The process initiates when a user submits a natural language request (Step 1). The frontend forwards this to the backend (Step 2), which performs intent recognition. If a modeling task is identified, the backend queries the LLM Service (Step 3). Unlike standard chatbots, the LLM is instructed to return structured JSON data representing the process or specific editing commands (Step 4). The backend then converts this JSON into standard BPMN 2.0 XML and transmits it to the Layout Server (Step 5). The layout server calculates the X/Y coordinates for the diagram elements and returns the enriched XML (Step 6). Finally, the backend delivers the renderable XML to the frontend for visualization (Step 7).
@@ 3.5 JSON Representation of BPMN Diagrams
The JSON intermediate representation is not merely an implementation convenience, but a conceptual abstraction that captures BPMN control-flow semantics independently of any concrete serialization format. BPMN Assistant utilizes a hierarchical JSON representation to abstract the verbose XML syntax of standard BPMN 2.0. This structure is designed to be easily generated by LLMs while maintaining sufficient fidelity to map back to valid XML. The representation uses a sequence of elements to describe the process flow. Unless branching logic is introduced via gateways, elements in the process array are executed sequentially. By elevating process structure to a first-class representation, this abstraction allows LLMs to reason about process logic directly, without being exposed to the syntactic complexity of BPMN XML.
@@
Tasks represent atomic units of work. The system supports a comprehensive set of task types to handle various modeling scenarios. Beyond the generic task, the system supports: userTask, serviceTask, sendTask, receiveTask, businessRuleTask, manualTask, and scriptTask.
    "type": "userTask", // or serviceTask, sendTask, etc.
    "id": "task_123",
    "label": "Approve request"
@@
The system distinguishes between start, end, and intermediate events. Crucially, it supports eventDefinition attributes to define specific triggers, such as timers or messages. This allows for the modeling of event-driven architectures.
    "type": "intermediateCatchEvent",
    "id": "event_timer",
    "label": "Wait 24 hours",
    "eventDefinition": "timerEventDefinition"
@@
Gateways manage flow divergence and convergence. The system supports exclusive (XOR), inclusive (OR), and parallel gateways.
A unique feature of our representation is the has_join boolean attribute. Since the JSON structure is hierarchical (nested), has_join:true explicitly signals the XML transformer that the branches merge back into a single flow after execution, triggering the generation of a converging gateway node. If false, the branches may end independently or loop back to previous elements.
To support cyclic flows (loops) within a nested structure, branches utilize an optional next field. This field contains the ID of a target element, allowing the flow to jump to any existing node in the process, breaking the strict hierarchy when necessary.
    "type": "exclusiveGateway", // or "inclusiveGateway"
    "id": "gateway_1",
    "label": "Is approved?",
    "has_join": true,
@@
            "condition": "Yes",
            "path": [ ... ] // Nested sequence of elements
            "condition": "No",
            "path": [],
            "next": "task_start" // Example of a loop-back
For inclusive gateways, an is_default boolean field is available to designate the default flow path.
Parallel gateways are represented as an array of arrays, where each sub-array constitutes a concurrent execution path. Synchronization is handled implicitly: if the parent flow continues after the gateway, a converging parallel gateway is automatically generated in the XML output.
@@
        [ { "type": "task", "label": "Path A" } ],
        [ { "type": "task", "label": "Path B" } ]
@@ 3.6 Process Editing Functions
The system supports a set of specialized functions for modifying BPMN diagrams. These functions, outlined in Table 3, enable precise control over process elements while maintaining the structural integrity of the diagram. From a conceptual standpoint, these editing functions define the minimal set of operations required to express meaningful BPMN model transformations while preserving process soundness. When the LLM receives an editing request, it analyzes the natural language input and determines which function(s) to call to achieve the desired modifications.
@@
3.7 Validation and Soundness
To ensure the generation of syntactically sound BPMN models, the system incorporates a strict validation layer implemented in Python that intercepts the LLM output prior to XML conversion. This validator enforces core BPMN structural constraints, including the uniqueness of element identifiers, the correctness of connectivity between flow elements, the validity of gateway branch hierarchies, and the requirement that each process contains exactly one start event. When a violation is detected, the system initiates a self-correction loop in which the validation error is fed back to the LLM, prompting it to revise the intermediate representation. By acting as a programmatic guardrail, this validation mechanism ensures that the final BPMN XML is both syntactically valid and reliably interpretable by standard BPMN engines.
@@ 4 Evaluation
Our evaluation methodology deliberately focuses on structural and syntactic assessment through graph similarity measures. This approach was selected based on the research objectives of comparing intermediate representations for BPMN generation and modification. It is important to note that this evaluation does not encompass semantic evaluation approaches, such as those based on Petri nets or conformance checking from process mining. Similarly, we do not focus on execution performance through simulation studies, nor do we conduct comprehensive usability evaluations. While these aspects represent valuable dimensions for assessing process modeling tools, they fall outside the scope of the current research, which primarily examines the efficacy of different representation approaches in generating structurally accurate BPMN diagrams from natural language descriptions. The selected evaluation metrics allow for direct comparison between JSON and XML approaches while providing insight into the structural correctness of the generated models.
4.1 Dataset Composition
To ensure a robust evaluation across diverse scenarios, we constructed a dataset of 60 process descriptions spanning 20 distinct business domains (e.g., Marketing, Healthcare, Logistics). The dataset was curated to evaluate the system’s ability to handle fundamental BPMN control-flow constructs, including sequential tasks, exclusive decision points (XOR-split/join), and concurrent execution paths (AND-split/join). This focus on core structural patterns provides a baseline for assessing the generation reliability of the underlying JSON representation.
The evaluation set was generated using a fixed prompt template and the OpenAI gpt-4.1 model. Each description was constrained to 7–8 activities and formulated in natural language while explicitly discouraging BPMN-specific terminology. Three descriptions were generated for each of the 20 business domains, after which the outputs were manually screened to remove ambiguous or degenerate cases.
Ground-truth BPMN diagrams were then created from these descriptions by multiple BPMN-trained annotators, including the authors, following a consistent modeling guideline. All diagrams were subsequently reviewed and corrected by the authors to ensure syntactic validity and semantic alignment with the corresponding textual descriptions.